A tax auditor is selecting a sample of 5 tax returns for an audit. If 2 or more of these returns are​ "improper," the entire population of 40 tax returns will be audited. Complete parts​ (a) through​ (e) below.

a. What is the probability that the entire population will be audited if the true number of improper returns in the population is 15​?

The probability is ___?

Include Excel formula.
Answer #1

a) P(Entire population will be audited)

= P(2 or more improper returns in the sample of 5 when population is 40 and total improper returns are 15)

= 1 - P(Less than 2 improper returns)

= 1 - P(0 or 1 improper returns)

= 1 - (combin(25,5) + combin(15,1)*combin(25,4))/(combin(40,5)) [Excel Formula]

= 0.6309